OpenStudy (anonymous):

Write the quadratic equation in the variable x having the given numbers as solutions. the equation is standard form, ax^2+bx+c=0

OpenStudy (anonymous):

solution:5, only solution

OpenStudy (anonymous):

x=\[(-b \pm \sqrt{b^2-4ac})/(2a)\]

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

\[x ^{2}-10x+25\]=0

OpenStudy (anonymous):

solution:6, only solution

OpenStudy (anonymous):

\[x ^{2}-12x+36=0\]

OpenStudy (anonymous):

\[(x-6)^{2}=0\]

OpenStudy (anonymous):

use one since same...x-6=0 add 6 to both sides and get x=6